# Proper scoring rules for estimation and forecast evaluation

### v2

@@ 1 Introduction
where $F_{\mathbb{P}}$ is the cumulative distribution function (CDF) of $\mathbb{P}$ (Matheson and Winkler, 1976). It is a strictly proper scoring rule. For details, see Example 7.
@@
Proper scoring rules have long been recognized as elicitation and aggregation mechanisms for subjective opinion in economics and other social sciences (Jensen and Peterson, 1973). In fact, many interesting developments such as the spherical scoring rule originate there (Roby, 1964). Tetlock (2005) provides a popular account of studies on the accuracy of probabilistic forecasts concerning political events by subject-matter experts. Another interesting application is the crowd-sourcing of forecasts from multiple human subjects via forecasting tournaments such as the Good Judgement Open and Metaculus, which evaluate forecasts using the Brier and logarithmic scores, respectively (GJOpen, 2025; Metaculus, 2025). A fascinating analysis of the literature on proper scoring rules in applications from 1950 to 2015 is given by Carvalho (2016).
The literature knows both orientations in that a proper scoring rule $S(\mathbb{P},\mathbb{Q})$ should be minimized at $\mathbb{P}=\mathbb{Q}$ or that it should be maximized. The influential paper by Gneiting and Raftery (2007) assumes that better forecasts correspond to higher scoring rule values, but most of the more recent literature has adopted the convention in Equation 1. With a negative orientation, scoring rules should maybe be called loss functions, but it is not in our power to change the naming convention.
Some parts of the literature order the arguments of proper scoring rules differently, that is, it would be $S(y,\mathbb{P})$ instead of $S(\mathbb{P},y)$. We mention this since it may cause confusion when dealing with expected scores $S(\mathbb{P},\mathbb{Q})$. “Scoring rule” is sometimes abbreviated to “score”, which may cause confusion, since “score” typically refers to the gradient of the logarithmic density in statistics, machine learning and econometrics. However, for some classes of proper scoring rules such as kernel scores (Section 3.1) the term is established by now.
@@ 1.1 Proper scoring rules in estimation
which yields the estimator $\smash{\mathbb{P}_{\hat{\theta},x}}$ of $\mathbb{P}_{Y|X=x}$ for every $x\in\mathcal{X}$. This approach estimates the conditional distribution simultaneously for every quantile and covariate as opposed to some classical methods such as quantile regression. In this sense, strictly proper scoring rules play the same role for estimating conditional distributions $\mathbb{P}_{Y|X=x}$, that the squared error loss plays for the conditional expectation $\mathbb{E}[Y|X=x]$. Moreover, in a certain sense, these are the only loss functions which can consistently estimate conditional distributions (see Theorem 8).
@@ 1.2 Proper scoring rules in forecast evaluation
While Equation 1 ensures that the optimal forecaster (or true model) will dominate all other forecasters (models or parameters) with regards to the comparison in Equation 3 in the long term, different proper scoring rules will rank competing imperfect forecasters differently even for infinite amounts of data, see for example Patton (2020). While this may seem as a disadvantage of ranking forecasts with respect to proper scoring rules, it is neither surprising nor is there any alternative. Therefore, the scoring rule must be chosen carefully for the comparison of empirical scores to be relevant to the application at hand, since it determines when one imperfect probabilistic forecast should be preferred over another. As a result, there are numerous proposals to tailor proper scoring rules such that they are discriminative with respect to certain regions of interest of the predictive distributions and outcomes, see Section 5.
@@ 2.1 Definition
Let $\mathcal{Y}=\mathbb{R}$ and $\mathcal{P}$ be the set of distributions with finite mean. The continuous ranked probability score $\operatorname{CRPS}:\mathcal{P}\to\overline{\mathbb{R}}$ given by $\operatorname{CRPS}(\mathbb{P},y)=\int\left(F_{\mathbb{P}}(x)-\mathbf{1}\{y\leq x\}\right)^{2}\,\mathrm{d}x$, where $F_{\mathbb{P}}$ is the CDF of $\mathbb{P}$, is a strictly proper scoring rule (Matheson and Winkler, 1976). It has compelling interpretation as the $L^{2}$ distance between the CDFs of $\mathbb{P}$ and the forecast $y$. Alternatively, it can also be seen as the Brier score for the binary event $\mathbf{1}{\{Y\leq x\}}$ integrated over $\mathbb{R}$. Furthermore, it has the equivalent representation
@@
see Remark 1 in the appendix for derivation. The associated divergence $d(\mathbb{P},\mathbb{Q})=\tfrac{1}{2}\int\|\nabla\log p(\mathbf{y})-\nabla\log q(\mathbf{y})\|^{2}\,q(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is sometimes called score matching distance.
@@
and a symmetric divergence $d(\mathbb{P},\mathbb{Q})=\int(F_{\mathbb{P}}(y)-F_{\mathbb{Q}}(y))^{2}\,\mathrm{d}y$, whose square root is a metric, compare Section 3.1.
@@ 2.2 Characterization
for all $\mathbb{Q}\in\mathcal{P}$. The function $H$ in Equation 7 indeed coincides with the entropy of $S$ defined in Equation 6. A regular proper scoring rule is thus determined by its entropy $H$ and the family $\{h_{\mathbb{P}}:\mathbb{P}\in\mathcal{P}\}$ of chosen supergradients. Theorem 12 is general but also somewhat abstract since supergradients of concave functions on convex sets of probability measures may not be directly intuitive. However, we will showcase the strength and direct applicability of this result in several instances.
@@
for every $\mathbb{Q}\in\mathcal{P}$, and $\nabla_{\mathbb{P}}H$ is called gradient of $H$. If $H$ is differentiable, then supergradients are unique and equal to the gradient $\nabla_{\mathbb{P}}H=h_{\mathbb{P}}$ (see Ekeland and Témam, 1999, Proposition 5.3). Thus, for differentiable entropies, there is a bijective correspondence between a scoring rule and its entropy and Equation (7) provides a generic recipe for constructing proper scoring rules out of differentiable concave functions on $\mathcal{P}$.
@@ 2.3 Geometric Properties
Generally, the divergence of a proper scoring rule does not correspond to a metric but behaves like a (squared) metric in many respects. Locally speaking, the divergence can be said to impose a manifold structure on the space of probability measures with the Hessian $\nabla_{\mathbb{P}}^{2}H$ of the entropy acting as the metric tensor by relating the “distance” $d(\mathbb{P}+\Delta\mathbb{P},\mathbb{P})$ between two infinitesimally close points $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ to the difference $\Delta\mathbb{P}=\mathbb{Q}-\mathbb{P}$ between their “coordinates” $\mathbb{P}$ and $\mathbb{P}+\Delta\mathbb{P}$ as
$d(\mathbb{P},\mathbb{P}+\Delta\mathbb{P})\approx\frac{1}{4}\langle\Delta\mathbb{P},[\nabla_{\mathbb{P}}^{2}H]\Delta\mathbb{P}\rangle=\frac{1}{4}\int\nabla_{\mathbb{P}}^{2}H(x,x^{\prime})\,\mathrm{d}\Delta\mathbb{P}(x)\,\mathrm{d}\Delta\mathbb{P}(x^{\prime}),$ (10)
@@
$d(\mathbb{P},\mathbb{Q})=\frac{1}{2}\iint\left[\int_{0}^{1}t\nabla_{t\mathbb{P}+(1-t)\mathbb{Q}}^{2}H(x,x^{\prime})\,\mathrm{d}t\right]\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x)\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x^{\prime})$
@@
$\frac{1}{4}\int(q(\mathbf{x})/p(\mathbf{x})-1)^{2}\,\mathrm{d}\mathbf{x},\;\frac{1}{2}\int(q(\mathbf{x})-p(\mathbf{x}))^{2}\,\mathrm{d}\mathbf{x},\;-\frac{1}{8}\iint|x-y|(q(\mathbf{x})-p(\mathbf{x}))(q(\mathbf{y})-p(\mathbf{y}))\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}.$
The divergence also satisfies a generalization of the law of cosines (and Pythagoras theorem). Indeed, the so-called three point property holds, which means that for $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3}\in\mathcal{P}$,
@@ 3.1.2 Characterizations
The divergence of a proper scoring rule is generally not symmetric, although locally it behaves like a squared metric such as discussed in Section 2.3. Interestingly, the square root of the divergence of a kernel score is always a valid metric on $\mathcal{P}$. In fact, under mild conditions, these are the only proper scoring rules with this property.
@@
Let $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ be a proper scoring rule such that $\{\delta_{y}:y\in\mathcal{Y}\}\subseteq\mathcal{P}$ and the supergradient map $\mathbb{P}\mapsto h_{\mathbb{P}}$ is weakly continuous. Then, the divergence $d$ of $S$ is symmetric if and only if $S$ is strongly equivalent to a kernel score.
@@
Further important notions of equivariance and invariance are homogeneity and isotropy. A proper scoring rule $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ is said to be homogeneous of degree $\alpha$ if $S(\mathbb{P}_{c},c\mathbf{y})=c^{\alpha}S(\mathbb{P},\mathbf{y})$ for every $c>0$, $\mathbb{P}\in\mathcal{P}$ and $\mathbf{y}\in\mathbb{R}^{d}$ where $\mathbb{P}_{c}(A)=\mathbb{P}(c^{-1}A)$ for Borel sets $A\subseteq\mathbb{R}^{d}$. It is said to be isotropic or rotation-invariant if $S(\mathbb{P}_{\mathbf{U}},\mathbf{U}\mathbf{y})=S(\mathbb{P},\mathbf{y})$ for every rotation matrix $\mathbf{U}\in\mathrm{SO}(d)$, $\mathbb{P}\in\mathcal{P}$ and $\mathbf{y}\in\mathbb{R}^{d}$ where $\mathbb{P}_{\mathbf{U}}(A)=\mathbb{P}(\mathbf{U}^{\top}A)$ for Borel sets $A\subseteq\mathbb{R}^{d}$. A consequence of homogeneity, that is of great practical importance is that scaling the sample $\mathbf{y}$ by $c>0$ as $c\mathbf{y}$, scales the predictive distribution accordingly. Thus, using different units of measurement for the data (e.g. km/h and knots for wind speed) leads to the same optimal distribution in the corresponding units (in estimation contexts), or the same predictive performance assessment (in forecast evaluation contexts). The following result characterizes translation invariant kernel scores with these properties.
@@
Let $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ be a translation invariant kernel score with the matrix $\mathbf{B}$ and the spectral measure $\mu$ as in Theorem 20. Then there exists a $\sigma$-finite measure $\rho$ on $(0,\infty)$ with $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$ and probability measures $\{\nu_{r}:r\in(0,\infty)\}$ on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ such that
@@
    for some probability measure $\nu$ on $\mathbb{S}^{d-1}$ and the Lebesgue measure $\mathrm{d}r$ on $(0,\infty)$.
@@
In particular, the CRPS is the only $1$-homogeneous translation invariant kernel score on $\mathbb{R}$ and the energy scores are the only homogeneous isotropic translation invariant kernel scores on $\mathbb{R}^{d}$ (up to positive multiplicative constants).
@@ 3.1.3 Continuous ranked probability score (CRPS)
The CRPS is arguably the most popular kernel score for univariate data. It was originally derived as the limit of discrete ranked probability scores and by integrating the Brier score, until it was recognized as a member of a much larger family of kernel scores (Staël von Holstein, 1977; Gneiting and Raftery, 2007). Mathematically, it is essentially the only $1$-homogeneous translation invariant kernel score on $\mathbb{R}$ (see Theorem 21). In addition to historical reasons, its popularity is explained by a range of advantages that it offers.
The CRPS is an intuitive, interpretable, and convenient measure of forecasting accuracy. It is distance-sensitive, that is, it assigns better scores to forecasts, which put more probability closer to the observations. In contrast, the logarithmic score assigns harsh penalties regardless of the overall closeness of the forecast to the observations if the forecast density at any of the observations is too low. The CRPS is essentially the $L^{2}$ distance between the CDFs of the forecast and the observations (see Example 11), and as the integral of the quadratic (Brier) score for binary outcomes, it elegantly connects to the conceptually simpler point forecasting setting. Indeed, it reduces to the absolute error loss when point forecasts are issued. Moreover, homogeneity implies that scaling the observations scales the predictive distribution accordingly even when the CRPS is away from the minimum. Furthermore, the CRPS exists for all forecasts with finite first moment, which means that the forecast measure need not necessarily have a density. It can be easily estimated by Equation 14 using samples drawn from the forecast, or explicitly computed if the forecast distribution has finite support. This is especially useful in applications such as ensemble forecasting in meteorology, where forecasts are routinely issued as discrete predictive distributions consisting of the empirical distributions of the ensemble of predictions.

### v3

@@ 1 Introduction
where $F_{\mathbb{P}}$ is the cumulative distribution function (CDF) of $\mathbb{P}$ (Matheson and Winkler, 1976). It is a strictly proper scoring rule. See Example 7 for details.
@@
Proper scoring rules have long been recognized as elicitation and aggregation mechanisms for subjective opinion in economics and other social sciences (Jensen and Peterson, 1973). For example, the spherical scoring originates in psychology (Roby, 1964). Tetlock (2005) provides a popular account of studies on the accuracy of probabilistic forecasts concerning political events by subject-matter experts. Another interesting application is the crowd-sourcing of forecasts from multiple human subjects via forecasting tournaments such as the Good Judgement Open and Metaculus, which evaluate forecasts using the Brier and logarithmic scores, respectively (GJOpen, 2025; Metaculus, 2025). A fascinating analysis of the literature on proper scoring rules in applications from 1950 to 2015 is given by Carvalho (2016).
The extant literature contains both of the possible conventions for proper scoring rules: one which requires $S(\mathbb{P},\mathbb{Q})$ to be minimized at $\mathbb{P}=\mathbb{Q}$ and the other requiring it to be maximized instead. The influential paper by Gneiting and Raftery (2007) assumes that better forecasts correspond to higher scoring rule values, but most of the more recent literature has adopted the convention in Equation 1. With a negative orientation, scoring rules should be called loss functions, but it is not in our power to change the naming convention.
Some parts of the literature order the arguments of proper scoring rules differently, that is, it would be $S(y,\mathbb{P})$ instead of $S(\mathbb{P},y)$. We mention this since it may cause confusion when dealing with expected scores $S(\mathbb{P},\mathbb{Q})$. “Scoring rule” is sometimes abbreviated to “score”, which may cause confusion, since “score” typically refers to the gradient of the logarithmic density in statistics, machine learning and econometrics. However, for some classes of proper scoring rules such as kernel scores (Section 3.1) the term has now become established.
@@ 1.1 Proper scoring rules in estimation
which yields the estimator $\smash{\mathbb{P}_{\hat{\theta},x}}$ of $\mathbb{P}_{Y|X=x}$ for every $x\in\mathcal{X}$. This approach estimates the conditional distribution simultaneously for every quantile and covariate, unlike some classical methods such as quantile regression. In this sense, strictly proper scoring rules play the same role for estimating conditional distributions $\mathbb{P}_{Y|X=x}$, that the squared error loss plays for the conditional expectation $\mathbb{E}[Y|X=x]$. Moreover, in a certain sense, these are the only loss functions which can consistently estimate conditional distributions (see Theorem 8).
@@ 1.2 Proper scoring rules in forecast evaluation
While Equation 1 ensures that the optimal forecaster (or true model) will dominate all other forecasters (models or parameters) with regards to the comparison in Equation 3 in the long term, different proper scoring rules will rank competing imperfect forecasters differently even for infinite amounts of data, see for example Patton (2020). While this may seem as a disadvantage of ranking forecasts with respect to proper scoring rules, it is not surprising and there there is no alternative. Therefore, the scoring rule must be chosen carefully for the comparison of empirical scores to be relevant to the application at hand, since it determines when one imperfect probabilistic forecast should be preferred over another. As a result, there are numerous proposals to tailor proper scoring rules such that they are discriminative with respect to certain regions of interest of the predictive distributions and outcomes, see Section 5.
@@ 2.1 Definition
Let $\mathcal{Y}=\mathbb{R}$ and $\mathcal{P}$ be the set of distributions with finite mean. The continuous ranked probability score $\operatorname{CRPS}:\mathcal{P}\to\overline{\mathbb{R}}$ given by $\operatorname{CRPS}(\mathbb{P},y)=\int\left(F_{\mathbb{P}}(x)-\mathbf{1}\{y\leq x\}\right)^{2}\,\mathrm{d}x$, where $F_{\mathbb{P}}$ is the CDF of $\mathbb{P}$, is a strictly proper scoring rule (Matheson and Winkler, 1976). It has a compelling interpretation as the $L^{2}$ distance between the CDFs of $\mathbb{P}$ and the forecast $y$. Alternatively, it can also be seen as the Brier score for the binary event $\mathbf{1}{\{Y\leq x\}}$ integrated over $\mathbb{R}$. Furthermore, it has the equivalent representation
@@
see Remark 1 in the appendix for a derivation. The associated divergence $d(\mathbb{P},\mathbb{Q})=\tfrac{1}{2}\int\|\nabla\log p(\mathbf{y})-\nabla\log q(\mathbf{y})\|^{2}\,q(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is sometimes called score matching distance.
@@
and a symmetric divergence $d(\mathbb{P},\mathbb{Q})=\int(F_{\mathbb{P}}(y)-F_{\mathbb{Q}}(y))^{2}\,\mathrm{d}y$, whose square root is a metric, compare Example 17.
@@ 2.2 Characterization
for all $\mathbb{Q}\in\mathcal{P}$. The function $H$ in Equation 7 indeed coincides with the entropy of $S$ defined in Equation 6. A regular proper scoring rule is thus determined by its entropy $H$ and the family $\{h_{\mathbb{P}}:\mathbb{P}\in\mathcal{P}\}$ of chosen supergradients. Theorem 12 is general but also somewhat abstract since supergradients of concave functions on convex sets of probability measures may not be directly intuitive. However, in the following, we will showcase the strength and direct applicability of this result in several instances.
@@
for every $\mathbb{Q}\in\mathcal{P}$, and $\nabla_{\mathbb{P}}H$ is called the gradient of $H$. If $H$ is differentiable, then supergradients are unique and equal to the gradient $\nabla_{\mathbb{P}}H=h_{\mathbb{P}}$ (see Ekeland and Témam, 1999, Proposition 5.3). Thus, for differentiable entropies, there is a bijective correspondence between a scoring rule and its entropy and Equation (7) provides a generic recipe for constructing proper scoring rules out of differentiable concave functions on $\mathcal{P}$.
@@ 2.3 Geometric Properties
Generally, the divergence of a proper scoring rule does not correspond to a metric but behaves like a (squared) metric in many respects. Locally speaking, the divergence can be said to impose a manifold structure on the space of probability measures with the Hessian $\nabla_{\mathbb{P}}^{2}H$ of the entropy acting as the metric tensor by relating the “distance” $d(\mathbb{P},\mathbb{P}+\Delta\mathbb{P})$ between two infinitesimally close points $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ to the difference $\Delta\mathbb{P}=\mathbb{Q}-\mathbb{P}$ between their “coordinates” $\mathbb{P}$ and $\mathbb{P}+\Delta\mathbb{P}$ as
$d(\mathbb{P},\mathbb{P}+\Delta\mathbb{P})\approx-\frac{1}{4}\langle\Delta\mathbb{P},[\nabla_{\mathbb{P}}^{2}H]\Delta\mathbb{P}\rangle=-\frac{1}{4}\int\nabla_{\mathbb{P}}^{2}H(x,x^{\prime})\,\mathrm{d}\Delta\mathbb{P}(x)\,\mathrm{d}\Delta\mathbb{P}(x^{\prime}),$ (10)
@@
$d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint\left[\int_{0}^{1}t\nabla_{t\mathbb{P}+(1-t)\mathbb{Q}}^{2}H(x,x^{\prime})\,\mathrm{d}t\right]\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x)\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x^{\prime})$
@@
$\frac{1}{4}\int(q(\mathbf{x})/p(\mathbf{x})-1)^{2}p(\textbf{x})\,\mathrm{d}\mathbf{x},\;\frac{1}{2}\int(q(\mathbf{x})-p(\mathbf{x}))^{2}\,\mathrm{d}\mathbf{x},\;-\frac{1}{8}\iint|x-y|(q(\mathbf{x})-p(\mathbf{x}))(q(\mathbf{y})-p(\mathbf{y}))\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mathbf{y}.$
The divergence also satisfies a generalization of the law of cosines (and the Pythagorean theorem). Indeed, the so-called three point property holds, which means that for $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3}\in\mathcal{P}$,
@@ 3.1 Kernel scores
Kernel scores form a flexible and powerful class of proper scoring rules that generalize many familiar examples. They are based on reproducing kernel Hilbert space (RKHS) methods and offer a principled way to compare probability distributions using kernel-based embeddings. In what follows, we detail their construction and computational aspects.
@@ 3.1.2 Characterizations
The divergence of a proper scoring rule is generally not symmetric, although locally it behaves like a squared metric, as discussed in Section 2.3. Interestingly, the square root of the divergence of a kernel score is always a valid metric on $\mathcal{P}$. In fact, under mild conditions, these are the only proper scoring rules with this property.
@@
Let $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ be a regular proper scoring rule such that $\{\delta_{y}:y\in\mathcal{Y}\}\subseteq\mathcal{P}$ and the supergradient map $\mathbb{P}\mapsto h_{\mathbb{P}}$ is weakly continuous. Then, the divergence $d$ of $S$ is symmetric if and only if $S$ is strongly equivalent to a kernel score.
@@
Further important notions of equivariance and invariance are homogeneity and isotropy. A proper scoring rule $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ is said to be homogeneous of degree $\alpha$ if $S(\mathbb{P}_{c},c\mathbf{y})=c^{\alpha}S(\mathbb{P},\mathbf{y})$ for every $c>0$, $\mathbb{P}\in\mathcal{P}$, and $\mathbf{y}\in\mathbb{R}^{d}$ where $\mathbb{P}_{c}(A)=\mathbb{P}(c^{-1}A)$ for Borel sets $A\subseteq\mathbb{R}^{d}$. It is said to be isotropic or rotation-invariant if $S(\mathbb{P}_{\mathbf{U}},\mathbf{U}\mathbf{y})=S(\mathbb{P},\mathbf{y})$ for every rotation matrix $\mathbf{U}\in\mathrm{SO}(d)$, $\mathbb{P}\in\mathcal{P}$, and $\mathbf{y}\in\mathbb{R}^{d}$ where $\mathbb{P}_{\mathbf{U}}(A)=\mathbb{P}(\mathbf{U}^{\top}A)$ for Borel sets $A\subseteq\mathbb{R}^{d}$. A consequence of homogeneity, that is of great practical importance is that scaling the sample $\mathbf{y}$ by $c>0$ as $c\mathbf{y}$, scales the predictive distribution accordingly. Thus, using different units of measurement for the data (e.g. km/h and knots for wind speed) leads to the same optimal distribution in the corresponding units (in estimation contexts), or the same predictive performance assessment (in forecast evaluation contexts). The following result characterizes translation invariant kernel scores with these properties.
@@
Let $S:\mathcal{P}\times\mathbb{R}^{d}\to\overline{\mathbb{R}}$ be a translation invariant kernel score with the matrix $\mathbf{B}$ and the spectral measure $\mu$ as in Theorem 20. Then there exists a $\sigma$-finite measure $\rho$ on (0,)(0$\infty$)( 0 , ∞ ) with $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$ and probability measures $\{\nu_{r}:r\in(0,\infty)\}$ on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ such that
@@
    for some probability measure $\nu$ on $\mathbb{S}^{d-1}$ and the Lebesgue measure $\mathrm{d}r$ on (0,)(0$\infty$)( 0 , ∞ ).
@@
In particular, the CRPS is the only 111-homogeneous translation invariant kernel score on $\mathbb{R}$ and the energy scores are the only homogeneous isotropic translation invariant kernel scores on $\mathbb{R}^{d}$ (up to positive multiplicative constants).
@@ 3.1.3 Continuous ranked probability score (CRPS)
The CRPS is arguably the most popular kernel score for univariate data. It was originally derived as the limit of discrete ranked probability scores and by integrating the Brier score, until it was recognized as a member of a much larger family of kernel scores (Staël von Holstein, 1977; Gneiting and Raftery, 2007). Mathematically, it is essentially the only 111-homogeneous translation invariant kernel score on $\mathbb{R}$ (see Theorem 21). In addition to historical reasons, its popularity is explained by a range of advantages that it offers.
The CRPS is an intuitive, interpretable, and convenient measure of forecasting accuracy. It is distance-sensitive, that is, it assigns better scores to forecasts, which put more probability closer to the observations. In contrast, the logarithmic score assigns harsh penalties regardless of the overall closeness of the forecast to the observations if the forecast density at any of the observations is too low. The CRPS is essentially the $L^{2}$ distance between the CDFs of the forecast and the observations (see Example 11), and as the integral of the quadratic (Brier) score for binary outcomes, it elegantly connects to the conceptually simpler point forecasting setting. Indeed, it reduces to the absolute error loss when point forecasts are issued. Moreover, homogeneity implies that scaling the observations scales the predictive distribution accordingly even when the CRPS is away from the minimum. Furthermore, the CRPS exists for all forecasts with a finite first moment, which means that the forecast measure need not necessarily have a density. It can be easily estimated by Equation 14 using samples drawn from the forecast, or explicitly computed if the forecast distribution has finite support. This is especially useful in applications such as ensemble forecasting in meteorology, where forecasts are routinely issued as discrete predictive distributions consisting of the empirical distributions of the ensemble of predictions.